Can someone explain how to solve this question in detail


. Draw the tripeptide GWA at pH 1.7, 14 - Calculate the molecular mass for GWA - Draw the titration curve for GWA - Calculate the pl for GWA Estimate the extinction coefficient for GWA and calculate the absorbance of a 1 g/L solution of GWA determined in a standard 1 cm cuvet Is this a meaningful absorbance for quantitation?
Answer #1

GWA consists of Alanine, Glycine ,Trypsine having a molecular mass of 332.36.

The isoelectric point od this sequence is 5.98 (approx).

The isoelectronic point or isoionic point is the pH at which the amino acid does not migrate in an electric field i.e, the pH at which the amino acid is neutral, i.e. the zwitterion form is dominant.

THere are three cases to consider

  • neutral side chains

These amino acids are characterised by two pKas : pKa1 and pKa2 for the carboxylic acid and the amine respectively. The isoelectronic point will be the average of, these two pKas, i.e.   pI = 1/2 (pKa1 + pKa2). For the simplest amino acid, glycine, pKa1= 2.34 and pKa2= 9.6, pI = 5.97.

  • acidic side chains

The pI will be at a lower pH because the acidic side chain introduces an "extra" negative charge. So the neutral form exists under more acidic conditions when the extra -ve has been neutralised. In aspartic acid , the neutral form is dominant between pH 1.88 and 3.65, pI is halfway between these two values, i.e.   pI = 1/2 (pKa1 + pKa3), so pI = 2.77.

  • basic side chains

The pI will be at a higher pH because the basic side chain introduces an "extra" positive charge. So the neutral form exists under more basic conditions when the extra +ve has been neutralised. For example, for histidine, the neutral form is dominant between pH 6.00 and 9.17, pI is halfway between these two values, i.e.  pI = 1/2 (pKa2 + pKa3), so pI = 7.59.

In a nutshell the most important factor is the overall bascicity or the acidity of the peptide chain that we have to consider. Then accordingly we have to calculate by means of the pKa's.

Absorbance(A) = e . c. l e= molar absorbtion coefficient , c= concentration , l = length of the cuvette.

The extinction coefficient obtained for the sequence is 5690 cm-1 M-1.

A = 5690 * 1 * 1 = 5690 .( at an unknown wavelength)

To get meaningful results,

• Abs260 > approximately twice Abs230 (A260/A230 Ratio=2)

• Abs260 = 0.1 or more (indicative of a high enough concentration, ie solution not too dilute) . The absorbances at the 230nm and 260 nm is required.